OpenStudy (anonymous):

Determine whether the curve has a tangent at x=1. If it does, give its slope. f(x) {x^2-2, x1

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} x^2-2 & \text{if} & x \leq 1 \\ 1.5x-2.5 & \text{if} & x >1 \end{array} \right.\]

OpenStudy (anonymous):

it is continuous at x = 1 because the limit from the left and from the right is 1

OpenStudy (anonymous):

but it does not have a slope there. the derivative of \[x^2-2\] is \[2x\] and at x = 1 you get 2

OpenStudy (anonymous):

whereas the derivative of a line is the slope, so the derivative of \[1.5x-2.5\] is 1.5

OpenStudy (anonymous):

and since \[2\neq 1.5\] not differentiable at 1

OpenStudy (anonymous):

(A)1.5 (B)2 (C)2.5 (D)3 (E)No Tangent So the answer would be E?